Midland Company buys tiles and prints different designs on them for souvenir and gift stores. It buys the tiles from a small company in Europe, so at all times it keeps on hand a stock equal to the tiles needed for three months’ sales. The tiles cost $1.00 each and must be paid for in cash. The company has 28,000 tiles in stock. Sales estimates, based on contracts received, are as follows for the next six months:

January

12,100

February

18,200

March

13,400

April

13,200

May

9,500

June

7,800

Required:

a & b. Estimate purchases (in units) and cash required to make purchases in January, February, and March
Answer #1

Answer- a&b-

MIDLAND COMPANY
MERCHANDISE PURCHASE BUDGET
FOR THE PERIOD ENDED MARCH 31
PARTICULARS JANUARY FEBRUARY MARCH
Estimated unit sales 12100 18200 13400
Add- Estimated sales for next three months 18200+13400+13200= 44800 13400+13200+9500= 36100 13200+9500+7800= 30500
Total inventory needs 56900 54300 43900
Less- Beginning inventory 28000 44800 36100
Units to be purchased (a) 28900 9500 7800
Unit cost (b) $1.00 $1.00 $1.00
Estimated purchase cost (c=a*b) $28900 $9500 $7800
Answer #2

44,800 from January becomes the beginning inventory for Febuary.